Operations Report – February 9th

Crew 221 Operations Report 09Feb2020

SOL: 7

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: on (at 6pm)

Hours run: 11:00

From what time last night: 8:00 pm

To what time this morning: 6:00 am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before the generator is run at night) 80%

Diesel Reading – 97%

Station Propane Reading – 58%

Ethanol Free Gasoline – gallons

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 250 gals

Static to Loft Pump used – yes

Water in GreenHab – 211.6 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: no

Hours: (before EVA) : 129.9

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging)

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: no

Hours: 136.1

Beginning charge: 100%

Ending charge:

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: Suit 10 battery replacement performed.

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment continues, now watered with 4l/day. watering of the plants carried out as per the procedure

Summary of ScienceDome operations: 1 incubator running, microbiology experiments going on and cultures are being grown in agar plates.

Summary of RAM operations: robot assembly is ongoing and expected to continue, antenna assembly for radio communication experiment continued today.

Summary of any observatory issues: nominal

Summary of health and safety issues: Two crew members have a sore throat. Another crew member hit their right knee with a rock during sol 6 EVA

Questions, concerns, and requests to Mission Support: the oven was reconfigured to work with propane as the previous configuration was for natural gas, and that was determined to be the cause of the smell reported. Everything is operational now.

Copyright © The Mars Society. All rights reserved. | Main Site